A 130.0 −mL buffer solution is 0.100 M in NH3 and 0.135 M in NH4Br.

What mass of HCl can this buffer neutralize before the pH falls below 9.00? Express the mass in grams to three significant figures.

If the same volume of the buffer were 0.265 M in NH3 and 0.400 M in NH4Br, what mass of HCl could be handled before the pH falls below 9.00? Express the mass in grams to three significant figures.
Answer #1

A)PK, of ammonia = -log K = -log(1.8x104) = 4.74 PK, of ammonia = 14.00-1pK = 14.00 -4.74 = 9.26 - - - - - - - - - - - - - -